OpenStudy (anonymous):

HELP ME PLEASE:D Part 1: Provide a system of TWO equations in slope-intercept form, with only one solution. Using complete sentences, explain why this system has one solution.

hero (hero):

x + y = 27 x - y = 13 Systems of Two Linear Equations are equations that can have either zero or one solution. If a system has zero solutions, it means that they have the same slope and the lines never intersect. If the system has one solution, it means that the lines intersect each other. Two lines can only intersect each other at one point. In this particular case, we have two lines x + y = 27 x - y = 13 In slope-intercept form y = mx + b, you would have: y = 27-x y = x - 13 Notice that for the first equation, the slope is -1. For the second equation, the slope is 1. Since the slopes are not the same, we know that these two lines will intersect at one point.

OpenStudy (anonymous):

sry got distracted tnk u so much

hero (hero):

yw